# Multi-Label Bayesian Active Learning with Inter-Label Relationships

### v1

@@ 3.3.1 Label-wise sampling
In multi-label scenarios, labels often exhibit asymmetric correlations, where the present of one label, $m$, is highly correlated with another label, $n$, but not vice versa. Hierarchical structures within labels are a common example of this kind of relationship. To illustrate the impact on performance, we can consider hierarchical data: when asymmetric correlations exist, the selection of root-node labels often overlaps with that of corresponding leaf-node labels, which reduces the representativeness of the selected root labels Nakano et al. (2020).
To address this, our strategy introduces a mechanism to modify label-wise selection. If correlation between labels exceeds a threshold, $\sigma$, defined here as the standard deviation of a two-tailed normal distribution, we consider the label pair to be asymmetrically correlated. In such cases, only the primary label in the correlation chain, $m$, is selected for label-wise sampling. The model then allocates the per-label query size for each label using pseudo labels to refine the sampling pool. This approach not only improves correlation-aware sampling but also mitigates imbalances by ensuring relatively even representation across selected samples.
@@ 3.3.2 Negative-correlated label sampling
In multi-label learning, ensuring the model respects the exclusivity of certain labels is crucial for achieving accurate predictions Huang et al. (2017). When mutually exclusive labels, such as those that should not logically co-occur, are predicted together, it is often an indication of model bias or misguided learning Huang and Kang (2021); Perales-González et al. (2020). This misalignment can reduce the model’s effectiveness. Therefore, as the second subset for concentrated sampling, we select instances with negatively correlated labels that are not expected to co-occur in the label space.
To formalize this, we consider a pair of labels as mutually exclusive when the negative correlation coefficient $NegA(m,n)$ exceeds a threshold of $2\sigma$, where $\sigma$ represents the standard deviation. Based on predictions with pseudo labels, our model selects a query size of samples from these instances to refine the unlabeled pool, guiding the model with a specific focus on avoiding negative correlations.
@@ 3.3.3 Hard-to-learn label sampling
The third subset our model uses to refine the unlabeled pool focuses on hard-to-learn samples. These samples are typically characterized by low confidence and low variability, indicating instances where the model has difficulty making accurate predictions. Such samples often contain ambiguous or noisy features or lie near decision boundaries Chang et al. (2017); Yang and Xu (2020). In this study, we set a confidence threshold of 0.5 to identify hard-to-learn samples. Specifically, if the pseudo label obtained through Eq. (10) for all classes of an instance falls below this threshold, our model classifies the sample as hard to learn. To improve performance on these challenging instances while maintaining diversity in the sampling process, our model dynamically adjusts the sample size using a polynomial decay function, enabling more focused learning on difficult cases over time.
3.4 Correlation-aware Querying
$S_{AB}(f_{L}(x),y^{n})$ $=\sum^{K}_{m=1}\hat{A}(m,n)*S_{BR}(f_{L}(x),y^{m})$ (12)
$\hat{A}(m,n)$ $=\text{norm}(A(m,n)),\quad\text{where }m\neq n$
$=\frac{A(m,n)}{\alpha\cdot\text{max}(\sum A(\cdot,n))}$ (13)
With the correlation-based sampling strategy described in section  3.3, our model obtains a refined unlabeled data pool. Then, our model calculate correlation-aware beta scoring for the selected samples, and use it to cluster the final samples for annotation. This process employs a computation method similar to the attention mechanism introduced in transformer models Vaswani et al. (2017). Using Eq. (12), we score each prediction, where $S_{AB}$ incorporates the influence of other labels’ scores through the attention coefficient $\hat{A}(m,n)$ as the final score, accounting the correlation. Additionally, we introduce $\alpha$, a normalization parameter set to 2, to prevent over-estimating correlated uncertainty while preserving the original significance of each label’s score. This approach allows our model consider the impact of neighboring labels on informativeness, and the refined unlabeled pool enhances computational efficiency and deepens the analysis of label correlations. Algorithm 1 details the procedure for one iteration of our framework.
@@ 3.4 Correlation-aware Querying
Input: Labeled pool: $L$; Unlabeled pool: $U$; Model: $\Theta^{E}={\theta_{1},...,\theta_{E}}$; Query size: $N$; Per-label query size: $M$; Hard to learn query size: $Z$.
@@
4:  Update correlation matrix $A(m,n)$ and $NegA(m,n)$ with $L$ via Eq. (8) and Eq. (9)
@@
6:  Select $M$ label-wise samples per label to $U^{*}$
7:  Select $N$ negative-correlated samples with $NegA$ to $U^{*}$
8:  Select $Z$ hard to samples to $U^{*}$
@@ 4 Experiments
We collected four benchmark multi-label text datasets to analyze the performance and robustness of our framework Kementchedjhieva and Chalkidis (2023). Those datasets include: RCV1 Lewis et al. (2004): A news articles dataset from Reuters; UKLEX Chalkidis and Søgaard (2022): A collection of legal documents sourced from various categories within UK law; EURLEX Chalkidis et al. (2021): A set of descriptors from European legal information thesaurus extracted from the European Union’s legal database; MIMIC3 Johnson et al. (2016): A set of de-identified health records for medical diagnosis. Since the data in UKLEX, EURLEX, and MIMIC3 have two levels of labels, to retain uniformity, our study used the coarse level of labels. Following the method by Charte et al.Charte et al. (2015), we used mean imbalance ratio (MeanIR) to create synthetic datasets with varying imbalance ratios based on the modified RCV1 dataset, reduce the label size of RCV1 to ten ($K=10$) by selecting the most frequently occurring labels, enabling an evaluation of the model’s performance across different degrees of imbalance. Table 1 and Table 2 offer a detailed summary of these four datasets. We also introduced a new metric, termed CorrAvg, defined as $\sum_{m=1}^{K}\sum_{n=1}^{K}A(m,n)/(K\times K),m\neq n$, to quantify the degree of the inter-correlation within the label set. We compared the performance of different query methods across datasets with different correlation levels.
Dataset #Document Train/Test #Vocab./ #Label #MeanIR Train/Test #CorrAvg Train/Test
RCV1 24891/6223 104619/102 402/197 0.137/0.137
UKLEX 20000/8500 63157/18 7/6 0.026/0.024
EURLEX 55000/5000 160211/21 16/15 0.131/0.147
MIMIC 29999/10000 137678/19 127/101 0.321/0.320
@@
Dataset #Document Train/Test #Vocab./ #Label #MeanIR Train/Test #CorrAvg Train/Test
RCV1-T10-5 1200/600 25254/10 5/10 0.133/0.138
RCV1-T10-10 1200/600 25289/10 10/10 0.135/0.138
RCV1-T10-20 1200/600 24170/10 20/10 0.137/0.138
RCV1-T10-50 1200/600 25280/10 50/10 0.142/0.138
@@
4.1 Implemetation
We used Neural-Classifier Liu et al. (2019), implemented in Pytorch Paszke et al. (2019), as the code base. In our study, we exployed three mainstream models, TextCNN Zhang and Wallace (2015), TextRNN Liu et al. (2016), and DistilBERT Sanh et al. (2019), as the backbone classifiers. To enhance efficiency and performance, we applied the cold start strategy Zhu et al. (2019) with random initialization at the beginning of each active learning iteration, a method known for its applicability to real-world scenarios Frankle and Carbin (2018). All experiments were conducted on a single RTX3090 GPU. Following the setting of Tan et al. Tan et al. (2024), the maximum sequence length for the text data was set to 256, with each training iteration consisting of 80 epochs. We implemented an early stopping criterion with the patience of 20 epochs to prevent the model from falling into local optima or overfitting Du et al. (2019); Ying (2019). AdamW was used as the optimizer Loshchilov and Hutter (2017), with the learning rate tailored for each model: 5e-2 for TextCNN and TextRNN, and 5e-5 for DistilBERT. The hard-to-learn query size was set to 300, while the per-label query size was set to 50 for RCV1 and 100 for other datasets, due to differences in label space size.
@@ 4.2 Baselines
To conduct a comparative performance analysis, we adopted five state-of-the-art MLAL methods as baselines, including random sampling. Each baseline uses the same query parameters and backbone classifier to maintain consistency across experiments. Specifically, MMC Yang et al. (2009) applies maximal confidence to selecting data that induces the largest reduction in expected model loss. AUDI Huang and Zhou (2013) explores uncertainty and diversity in both instance and label spaces through label ranking and threshold learning. ADAPTIVELi and Guo (2013) integrates max-margin prediction uncertainty with label cardinality inconsistency to assess the unified informativeness of multi-label instances. BESRA Tan et al. (2024) utilizes the beta scoring rules within an expected loss reduction framework to evaluate informativeness and employs vector representations to maintain diversity. CMAL Yu et al. (2020) leverages the global label correlation matrix and label space sparsity with the uncertainty to query the most informative example-label pairs.
@@ 4.3 Results
Figures 2 and 3 present the quantitative performance of our proposed framework. Figure 2 compares the performance of CRAB and baseline methods across four datasets using BERT, with additional results for other models is provided in Appendix 7. Following Tan et al. (2024), we obtained the predictive distribution by training five ensemble models independently, each initialized with the same parameters for every AL iteration. The micro-F1 results show that our proposed model, CRAB, has consistently outperforms other AL methods across different text domains and network structures. Additionally, CRAB demonstrates robust performance on datasets with varying degrees of correlation, particularly compared with BESRA, suggesting that our strategy effectively models correlation during data selection. To examine the robustness of the model on imbalanced datasets, we conducted comparative experiments on synthetic datasets, with the results shown in Figure 3. CRAB maintains superior performance across synthetic datasets with different MeanIR values, demonstrating its capability to handle imbalanced datasets.
To further investigate the effectiveness of our model, Figures 4 and 5 present a qualitative analysis of CRAB. Figure 4 shows the MeanIR of the selected samples across AL iterations. Since MeanIR indicates imbalance, with lower values reflecting a more even data distribution, we observe that CRAB demonstrates a more balanced sample selection compared to other baseline models, which underscores its capacity to address data imbalance effectively. Figure 5 illustrates the trend of two categories of data within the unlabeled data pool: hard-to-learn data and negatively correlated data. Unlike random selection, CRAB strategically selects data that enhances model learning, thereby reducing misclassification of negatively correlated data. Additionally, CRAB improves performance on hard-to-learn data, helping the model becomes more robust and accurate. This targeted data selection contributes to a more balanced and adaptive learning process, ultimately leading to improved generaliztion across diverse data types.
Figure 2: Averaged micro-F1 score on Bert, averaged results with 5 random seeds.
@@
4.4 Ablation study
We conducted four experiments to examine whether the structure of CRAB improves MLAL performance by considering correlation. Taking into consideration the asymmetrically correlated label relationships, CRAB selects only the initial label in the correlation chain for per-label selection, thus avoiding duplicate selection of correlated labels. Figure 6(a) compares performance with and without considering asymmetrical correlations on the MIMIC dataset. The results illustrate that CRAB demonstrates superior performance and is more effective in querying indicative samples than when treating all labels equally. Figure 6(b) shows the benefits of sampling conflicted labels, with performance improvements becoming more pronounced in the later training stages.
Figure 7(a) illustrates how the correlation attention impact MLAL accuracy. To assess performance without correlation in score evaluation, we removed the correlation attention in Eq. (12), with the results shown by the blue line. Evidently, when positive label correlations are incorporated, CRAB performs more consistently throughout the experiment, indicating that our strategy effectively models inter-label relationships to make more informative queries. Additionally, we compared the micro-F1 score and computation time of the random sampling and clustering-based sampling during the refined unlabeled pool sampling. As shown in Figure 7(b), random sampling performs almost identically to clustering-based sampling, suggesting it can server as a replacement during the refined unlabeled pool selection. Moreover, the querying time with random sampling decreases by 40% . These findings demonstrate that our method is effective, efficient, and robust.

### v2

@@ 3.3.1 Label-wise sampling
In multi-label scenarios, labels often exhibit asymmetric correlations, where the present of one label, $m$, is highly correlated with another label, $n$, but not vice versa. Hierarchical structures within labels are a common example of this kind of relationship. To illustrate the impact on performance, we can consider hierarchical data: when asymmetric correlations exist, the selection of root-node labels often inevitably overlaps with that of corresponding leaf-node labels, while rarely occurring independently, which reduces the representativeness of the selected root labels [Nakano et al., 2020].
To address this, our strategy introduces a new mechanism for the label-wise selection. If one label $n$ is highly dependent on the presence of another $m$, while the reverse is not necessarily true—otherwise, they would effectively be considered the same label in most cases—and their correlation exceeds a predefined threshold, $\sigma$, set as the standard deviation of a two-tailed normal distribution, we classify the label pair as asymmetrically correlated. To improve label-wise sampling, we then refine the label space of instances that contain both labels $m$ and $n$ by removing label $n$, ensuring that sampling prioritizes the most independent label, $m$. The model then performs evenly sampling based on the refined pseudo labels, ensuring a more balanced sampling pool.
@@ 3.3.2 Negative-correlated label sampling
In multi-label learning, it is essential for the model to respect the exclusivity of certain labels to ensure accurate predictions [Huang et al., 2017]. When mutually exclusive labels, such as those that should not logically co-occur, are predicted together, it is often an indication of model bias or misguided learning [Huang and Kang, 2021, Perales-González et al., 2020]. This misalignment can reduce model’s effectiveness. As the second subset for concentrated sampling, we select instances with negatively correlated pseudo labels that are not expected to co-occur in the label space.
To formalize this, we consider a pair of labels as mutually exclusive when the negative correlation coefficient $NegA(m,n)$ exceeds a predefined threshold, set as $2\sigma$, where $\sigma$ represents the standard deviation. Based on predictions with pseudo labels, our model selects samples with the predicted labels that are unlikely to co-occur, according to the negative correlation matrix, as those samples are at high risk of incorrect predictions. These selected samples are then added to the refined subset of the unlabeled pool, ensuring that the model better accounts for negative correlations.
@@ 3.3.3 Hard-to-learn label sampling
The third set, which our model uses to further expand the subset, consists of hard-to-learn samples. These samples are typically characterized by low confidence and low variability, indicating instances where the model has difficulty making accurate predictions. Such samples often contain ambiguous or noisy features or lie near decision boundaries [Chang et al., 2017, Yang and Xu, 2020]. In this study, we define samples without any predicted pseudo labels as hard-to-learn samples. Specifically, if the pseudo labels obtained through Eq. (11) for all classes of a given instance falls below the threshold, 0.5, the classifier cannot make any prediction, thus the sample is classified as hard to learn. To improve performance on these challenging instances while maintaining diversity in the sampling process, our model dynamically adjusts the sample size using a polynomial decay function, enabling more focused learning on difficult cases over time.
@@
Input: Labeled pool: $L$; Unlabeled pool: $U$; Model: $\Theta^{E}={\theta_{1},...,\theta_{E}}$; Query size: $N$; Per-label query size: $N$; Hard to learn query size: $Z$.
@@
4:  Update correlation matrix $A$ and $NegA$ with $L$ via Eq. (9) and Eq. (10)
@@
6:  Select and add $N$ label-wise samples per label to $U^{*}$
7:  Select and add $N$ negative-correlated samples with $NegA$ to $U^{*}$
8:  Select and add $Z$ hard samples to $U^{*}$
@@
3.4 Correlation-aware Querying
With the correlation-based sampling strategy described in section  3.3, our model obtains a refined subset of the original unlabeled data pool. Then, our model calculates correlation-aware beta scores for these samples in the selected subset, and use those scores to cluster the samples for acquisition. This score computation method is inspired by the attention mechanism introduced in transformer models [Vaswani et al., 2017], which is defined as follows.
$S_{AB}(f_{L}(x),y)$ $=\sum^{K}_{m=1}\hat{A}(m,:)S_{BR}^{m}(f_{L}(x),y^{m})$ (13)
$\hat{A}(m,n)$ $=\text{norm}(A(m,n)),\quad\text{where }m\neq n$
$=\frac{A(m,n)}{\gamma\cdot\text{max}(A(:,n))}$ (14)
Using Eq. (13), we score each prediction, where $S_{AB}$ incorporates the influence of other labels’ scores through the attention coefficient $\hat{A}$ as the final score, accounting the correlation. Additionally, we introduce $\gamma$, a normalization parameter set to 2, to prevent over-estimating correlated uncertainty while preserving the original significance of each label’s score. This approach allows our model consider the impact of neighboring labels on informativeness, and the refined unlabeled pool enhances computational efficiency and deepens the analysis of label correlations. Algorithm 1 details the procedure for one iteration of our framework.
@@ 4 Experiments
We collected four benchmark multi-label text datasets to analyze the performance and robustness of our framework [Kementchedjhieva and Chalkidis, 2023]. Those datasets include: RCV1 [Lewis et al., 2004], UKLEX [Chalkidis and Søgaard, 2022], EURLEX [Chalkidis et al., 2021], and MIMIC3 [Johnson et al., 2016]. RCV1 is a news articles dataset from Reuters; UKLEX is a collection of legal documents sourced from various categories within UK law; EURLEX is a set of descriptors from European legal information thesaurus extracted from the European Union’s legal database; and MIMIC3 is a set of de-identified health records for medical diagnosis. Following the method by Charte et al. [2015], we used mean imbalance ratio (MeanIR) to create synthetic datasets with varying imbalance ratios based on the modified RCV1 dataset, reduce the label size of RCV1 to ten by selecting the most frequently occurring labels, enabling an evaluation of the model’s performance across different degrees of imbalance. Table 1 and Table 2 offer a detailed summary of these four datasets. We also introduced a new metric, termed CorrAvg, defined as $\sum_{m=1}^{K}\sum_{n=1}^{K}$ $A(m,n)/(K\times K),m\neq n$, to quantify the degree of inter-correlation within label set.
4.1 Implemetation
We used Neural-Classifier [Liu et al., 2019], implemented in Pytorch [Paszke et al., 2019], as the code base. In our study, we exployed three mainstream models, TextCNN [Zhang and Wallace, 2017], TextRNN [Liu et al., 2016], and DistilBERT [Sanh, 2019], as the backbone classifiers. To enhance efficiency and performance, we applied the cold start strategy [Zhu et al., 2019] with random initialization at the beginning of each active learning iteration, a method known for its applicability to real-world scenarios [Frankle and Carbin, 2018]. All experiments were conducted on a single RTX3090 GPU. Following the setting of Tan et al. [2024], the maximum sequence length for the text data was set to 256, with each training iteration consisting of 80 epochs. The initial training set and validation set sizes are set to 100 and 1000, respectively, and are sampled from the training set. We implemented an early stopping criterion with the patience of 30 epochs to prevent the model from falling into local optima or overfitting [Du et al., 2019, Ying, 2019]. AdamW was used as the optimizer [Loshchilov and Hutter, 2019], with the learning rate tailored for each model: 5e-2 for TextCNN and TextRNN, and 5e-5 for DistilBERT. The hard-to-learn query size was set to 300 for benchmark datasets and 200 for synthetic datasets, while the per-label query size was set to 50 for RCV1 and 100 for other datasets, due to differences in label space size.
Dataset #Document Train/Test #Vocab./ #Label #MeanIR Train/Test #CorrAvg Train/Test RCV1 24,891/6223 104,619/102 402/197 0.137/0.137 UKLEX 20,000/8500 63,157/18 7/6 0.026/0.024 EURLEX 55,000/5000 160,211/21 16/15 0.131/0.147 MIMIC 29,999/10000 137,678/19 127/101 0.321/0.320
@@ 4.1 Implemetation
Dataset #Document Train/Test #Vocab./ #Label #MeanIR Train/Test #CorrAvg Train/Test RCV1-T10-5 1,200/600 25,254/10 5/10 0.133/0.138 RCV1-T10-10 1,200/600 25,289/10 10/10 0.135/0.138 RCV1-T10-20 1,200/600 24,170/10 20/10 0.137/0.138 RCV1-T10-50 1,200/600 25,280/10 50/10 0.142/0.138
@@ 4.2 Baselines
To conduct a comparative performance analysis, we adopted five state-of-the-art MLAL methods as baselines, including random sampling. Each baseline uses the same query parameters and backbone classifier to maintain consistency across experiments. Specifically, MMC [Yang et al., 2009] applies maximal confidence to selecting data that induces the largest reduction in expected model loss. AUDI [Huang and Zhou, 2013] explores uncertainty and diversity in both instance and label spaces through label ranking and threshold learning. ADAPTIVE[Li and Guo, 2013] integrates max-margin prediction uncertainty with label cardinality inconsistency to assess the unified informativeness of multi-label instances. BESRA [Tan et al., 2024] utilizes the beta scoring rules within an expected loss reduction framework to evaluate informativeness and employs vector representations to maintain diversity. CMAL [Yu et al., 2020] leverages global label correlation matrix and label space sparsity with uncertainty to query the most informative example-label pairs.
Figure 2: Averaged micro-F1 score on DistilBERT, averaged results with 5 random seeds.
@@ 4.3 Results
Figures 25 present the quantitative performance of our proposed framework. Figure 2 compares the performance of CRAB and baseline methods across four datasets using DistilBERT. Figure 3 and Figure 4 present supplementary performance comparisons based on TextCNN and TextRNN. Following Tan et al. [2024], we obtained the predictive distribution by training five ensemble models independently, each initialized with the same parameters for every AL iteration. The micro-F1 results show that our proposed model, CRAB, has consistently outperforms other AL methods across different text domains and network structures. Additionally, CRAB demonstrates robust performance on datasets with varying degrees of correlation, particularly compared with BESRA, suggesting that our strategy effectively models correlation during data selection.
Among the baseline methods, BESRA achieves strong results and shows relatively robustness on different datasets. However, its performance on the highly correlated MIMIC dataset is less stable, likely due to the absence of correlation consideration. AUDI, which incorporates both uncertainty and diversity at both data and instance level, presents notable performance on three of the datasets, RCV1, UKLEX, and EURLEX, but struggles on MIMIC. Adaptive and MMC yield similar results over four datasets, as both utilize the max-margin as the selection criterion. Although CMAL considers global label correlation, it only performs optimally on the highly correlated MIMIC dataset and does not maintain stable performance on all datasets.
Figure 3: Averaged micro-F1 score on TextCNN, averaged results with 5 random seeds.
To examine robustness of the model on imbalanced datasets, we conducted comparative experiments on synthetic datasets, with results shown in Figure 5. CRAB maintains superior performance across synthetic datasets with different MeanIR values, demonstrating its capability to handle imbalanced datasets.
To further investigate the effectiveness of our model, Figures 6 and 7 present a qualitative analysis of CRAB. Figure 6 shows the MeanIR of the selected samples across AL iterations. Since MeanIR indicates imbalance, with lower values reflecting a more even data distribution, we observe that CRAB demonstrates a more balanced sample selection compared to other baseline models, which underscores its capacity to address data imbalance effectively. Figure 7 illustrates the trend of two categories of data within the unlabeled data pool: hard-to-learn data and negatively correlated data. Unlike random selection, CRAB strategically selects data that enhances model learning, thereby reducing misclassification of negatively correlated data. Additionally, CRAB improves performance on hard-to-learn data, helping the model becomes more robust and accurate. This targeted data selection contributes to a more balanced and adaptive learning process, ultimately leading to improved generaliztion across diverse data types.
Figure 4: Averaged micro-F1 score on TextRNN, averaged results with 5 random seeds.
@@
4.4 Parameter sensitivity analysis
To validate the effectiveness and generalizability of CRAB, we conduct two experiments to analyze its performance under different parameter settings. Figure 8(a) demonstrates the performance with varying sizes of hard-to-learn samples for refined unlabeled pool selection. With an acquisition size of 100 per iteration, the model achieves the optimal performance. However, including any hard-to-learn samples in the sampling pool leads to a performance decline. If the sample size is set too large, such as 200, the model initially shows relatively better performance due to the higher proportion of hard-to-learn samples in the early stages. However, as annotated data increases, performance declines because the hard-to-learn samples become less influential, necessitating a reduction in their selection. This parameter is adjustable across different datasets and model structures to ensure compatibility with the learning capabilities across varying scenarios. To deal with the problem of the amount of hard-to-learn samples decreasing with increased annotated data, CRAB adopts a decay function for the size of hard-to-learn samples to adapt to the training process. Figure 8(b) presents performance of three decay approaches, linear decay, cosine decay, and polynomial decay. Among these, polynomial decay achieves superior performance in terms of micro-F1 score, as it produces an accelerated decrease in output for sampling, better aligning with the trend in the size of hard-to-learn samples.
@@ 4.4 Parameter sensitivity analysis
Figure 8: (a) Performance for different size of hard-to-learn samples. (b) Performance for different decay functions of the hard-to-learn samples.
4.5 Ablation study
We conducted four experiments to examine whether the structure of CRAB improves MLAL performance by considering correlation. Taking into consideration the asymmetrically correlated label relationships, CRAB selects only the initial label in the correlation chain for per-label selection, thus avoiding duplicate selection of correlated labels. Figure 9(a) compares performance with and without considering asymmetrical correlations on the MIMIC dataset. Results illustrate that CRAB demonstrates superior performance and is more effective in querying indicative samples than when treating all labels equally. Figure 9(b) shows the benefits of sampling conflicted labels, with performance improvements becoming more pronounced in later training stages.
Figure 10(a) illustrates how the correlation attention impact MLAL accuracy. To assess performance without correlation in score evaluation, we removed the correlation attention in Eq. (13), with the results shown by the blue line. Evidently, when positive label correlations are incorporated, CRAB performs more consistently throughout the experiment, indicating that our strategy effectively models inter-label relationships to make more informative queries. Additionally, we compared the micro-F1 score and computation time of the random sampling and clustering-based sampling during the refined unlabeled pool sampling. As shown in Figure 10(b), random sampling performs almost identically to clustering-based sampling, suggesting it can serve as a replacement during the refined unlabeled pool selection. Moreover, the querying time with random sampling decreases by 40% . These findings demonstrate that our method is effective, efficient, and robust.